12. An uncharged capacihor (capacitonce C) is connected in series of the followin 13. An inductor (inductance L.) is connected in a series circuit with a 9V battery, a resistor (resistance R) and a switch. The switch is closed at t-0. Which of the following describes the current in the circuit as a function of time? 9V four different sets of wires that cross each other without actually ll four cases, and the directions of
Answer #1

Current in the capacitor is given by

i = \frac{dq}{dt}

Here charge is given by

q = CE[1 - e^{-\frac{t}{RC}}]

now current is given by

i = \frac{dq}{dt}

i = \frac{d}{dt} (CE[1 - e^{-\frac{t}{RC}}])

i = \frac{E}{R}e^{-\frac{t}{RC}}

i = \frac{9V}{R}e^{-\frac{t}{RC}}

PART B)

Now for inductor we have

i = \frac{9V}{R}[1 - e^{-\frac{Rt}{L}}]

So option (d) is correct